From California to New York, legislative bodies across the United States are considering eliminating or reducing the surcharges that banks impose on noncustomers, who make $14 million in withdrawals from other banks’ ATM machines. On average, noncustomers earn a wage of $20 per hour and pay ATM fees of $3.25 per transaction. It is estimated that banks would be willing to maintain services for 5 million transactions at $1.25 per transaction, while noncustomers would attempt to conduct 22 million transactions at that price. Estimates suggest that, for every 1 million gap between the desired and available transactions, a typical consumer will have to spend an extra minute traveling to another machine to withdraw cash. Based on this information, what would be the nonpecuniary cost of legislation that would place a $1.25 cap on the fees banks can charge for noncustomer transactions? Instructions: Enter your responses rounded to the nearest penny (two decimal places). $ What would be the full economic price of this legislation? $
Answer #1

On average, noncustomers earn a wage of $20 per hour and pay ATM fees of $3.25 per transaction. This means:

  • Equilibrium price, Pe = $ 3.25 / transaction
  • Opportunity cost = $ 20 / hour

Cap on the fees banks can charge for non customer transactions = $ 1.25. This implies:

  • Ceiling price,C = $ 1.25 / transaction

t is estimated that banks would be willing to maintain services for 5 million transactions at $1.25 per transaction, while non customers would attempt to conduct 22 million transactions at that price. This implies:

  • Shortage in transaction volumes = 22 mn - 5 mn = 17 mn

For every 1 million gap between the desired and available transactions, a typical consumer will have to spend an extra minute traveling to another machine to withdraw cash. This implies:

  • Time lost in 17 mn of shortage transaction = 17 mn x 1 minute / mn = 17 minutes = 17 / 60 hours
  • Opportunity cost of 17 minutes = $ 20 / hour x 17 / 60 = $ 5.67

What would be the non pecuniary cost of legislation that would place a $1.25 cap on the fees banks can charge for non customer transactions?

Thus, the non pecuniary cost of legislation = opportunity cost = $ 5.67 per transaction

What would be the full economic price of this legislation?

Full economic price of this legislation = Ceiling price + opportunity cost = $ 1.25 + 5.67 = $ 6.92 / transaction